OpenStudy (anonymous):

Multiply these rational expressions. Help?:

OpenStudy (anonymous):

\[\frac{ 5x }{ 10x+5 } \times \frac{ 4x+2 }{ 5}\]

OpenStudy (anonymous):

5x(4x+2)/ 5(10x+5)

OpenStudy (anonymous):

Sorry: I meant: \[\frac{ 20x(x+2) }{ 50(x+5) }\]

OpenStudy (anonymous):

Do I do anything further with it?

OpenStudy (anonymous):

\[ \frac{5x}{10x + 5} \times \frac{4x + 2}{5} \] can be written as \[ \frac{5 (x) }{5 \left( 2x + 1 \right) } \times \frac{ 2 \left( 2x + 1 \right) }{5} \] Try to build on this.

OpenStudy (anonymous):

I do not believe your answer is correct

OpenStudy (anonymous):

if I am right, the answer should be \[ \frac{2x}{5} \]

OpenStudy (anonymous):

Okay, I see what I did wrong. Thank you!